### v1

@@ Abstract
Surprising videos, e.g., funny clips, creative performances, or visual illusions, attract significant attention. Enjoyment of these videos is not simply a response to visual stimuli; rather, it hinges on the human capacity to understand (and appreciate) commonsense violations depicted in these videos. We introduce FunQA, a challenging video question answering (QA) dataset specifically designed to evaluate and enhance the depth of video reasoning based on counter-intuitive and fun videos. Unlike most video QA benchmarks which focus on less surprising contexts, e.g., cooking or instructional videos, FunQA covers three previously unexplored types of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. For each subset, we establish rigorous QA tasks designed to assess the model’s capability in counter-intuitive timestamp localization, detailed video description, and reasoning around counter-intuitiveness. We also pose higher-level tasks, such as attributing a fitting and vivid title to the video, and scoring the video creativity. In total, the FunQA benchmark consists of 312K free-text QA pairs derived from 4.3K video clips, spanning a total of 24 video hours. Extensive experiments with existing VideoQA models reveal significant performance gaps for the FunQA videos across spatial-temporal reasoning, visual-centered reasoning, and free-text generation.
@@ 1 Introduction
The charm of surprising videos, be they funny, creative, or filled with visual illusions, offer enjoyment and commands attention from viewers. This type of media elicits positive surprise [1],111c.f., negative surprise, e.g., a surprising medical bill. a captivating emotion that stems not merely from perceiving surface-level visual stimuli, but rather, the innate ability of humans to understand and find delight in unexpected and counter-intuitive moments [2]. However, despite significant advancements in today’s computer vision models, the question remains: can video models “understand" the humor/creativity in surprising videos?
Consider the humorous video depicted in Figure 1 (left) as an illustrative example. We witness a man engrossed in his phone, sharing a meal with friends. Suddenly, one of his companions squeezes a generous amount of ketchup, which, instead of adorning the fries, splatters onto the man’s face. The shock in his eyes, combined with his ketchup-covered visage, elicits laughter. [^2] The hostility/superiority theory of humor posits that humor can arise from claiming superiority over someone or something [3, 4]; but alternate (more optimistic) theories of humor exist, [5] offers a survey. While humans effortlessly recognize this as an unusual (and potentially entertaining) event, the reasoning required to holistically understand the scene is complex: a model needs to recognize that individuals were gathered to enjoy a meal together, and discern that the comedic element arises from the ketchup intended for the fries ending up on the man’s face instead, and that the innocent expression of the unfortunate victim indicates no significant harm was caused.
@@
While there have been some efforts to enhance computer vision models’ performance in Video Question Answering (VideoQA), these works have primarily focused on the common, less surprising videos found in existing VideoQA datasets. Examples of commonly employed VideoQA datasets include YouCook2 [6] which contains video clips from 2K cooking videos, Howto100M [7] which consists of only instructional videos. While there exist video datasets that explore the humor in TV shows [8, 9] and include tasks such as predicting laughter tracks [10], these tasks often heavily rely on audio and narrative cues, with visual clues might playing a lesser role.
@@
To address this gap and evaluate computer vision models’ ability to identify and understand visual commonsense violations in videos, we introduce FunQA, a comprehensive and high-quality VideoQA dataset comprising 4.3K surprising videos and 312K manually annotated free-text QA pairs. Our dataset consists of three subsets: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset covers different sources and video contents, but the commonality lies in their surprising nature, e.g., the unexpected contrasts in humorous videos, the intriguing disguises in creative videos, and the seemingly impossible performances in magic videos. Our experiments suggest that these surprising videos require different types of reasoning than common videos, as existing VideoQA methods perform poorly on the corpus. With FunQA, we hope to provide a benchmark that covers the popular, important, and sophisticated genre of counter-intuitive/surprising videos.
@@
In FunQA, we formulate three rigorous tasks to measure models’ understanding of surprise: 1) Counter-intuitive timestamp localization: a model must identify the specific time period within a video when an unexpected event takes place. 2) Detailed video description: a model must generate coherent and objective descriptions of the video content, evaluating models’ fundamental video understanding capabilities. 3) Counter-intuitiveness reasoning: a model must generate concrete explanations of why the video is surprising. These tasks progressively assess the model’s ability to perceive, articulate, and reason about the counter-intuitive elements present in surprising videos. Additionally, we propose auxiliary tasks that pose higher-level challenges within the benchmark including assigning an appropriate and vivid title to the video, etc. To summarize our contributions:
@@
2) Novel and Challenging Tasks: We design a number of novel tasks that allow the model to explore previously untouched problems, such as timestamp localization, and reasoning around counter-intuitiveness. These tasks push video reasoning beyond superficial descriptions, demanding deeper understanding and discernment.
3) Comprehensive Evaluation: We have done an extensive and comprehensive evaluation of cutting-edge baselines, giving the field an insight and future research direction.
@@ 2 Related Work
Table 1: Comparison between FunQA and other existing benchmarks. Compare to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. Anno.: Annotation; M: Manual, A: Automatic; For Input, V, A, S, and B denote Video, Audio, Subtitle, and Bounding-box; VC means visual-centric, Desc. means Description, Expl for Explanation, STR for Spatial-temporal Reasoning. For QA Tasks, MC denotes Multiple Choice QA, OE means Open Ended QA, and FT means Free Text QA.
Video Question Answer
Datasets Domain Anno. Avg length (s) # Clips (K) Input # QA pairs (K) VC Desc. Expl. STR QA Task
MarioQA [11] Games M 3-6 188 V, A 188 Yes No Yes Yes OE
TGIF-QA [12] Social Media M 3.1 71.7 V 165.2 Yes Yes No Yes MC & OE
MovieQA [13] Movies A 202.7 6.77 V, A, S 6.4 No No Yes Yes MC
CLEVRER [14] Synthetic Video M 5 20 V 305 Yes No Yes Yes OE
TVQA [15] TV shows M 76.2 21.8 V, A 152.5 No No Yes No MC
TVQA+ [16] TV shows M 7.2 4.2 V, A, B 29.4 No No Yes Yes OE
Social-IQ [17] Web videos M 99 1.25 V, A, S, B 7.5 Yes No Yes No MC
NExT-QA [18] Daily life M 44 5.4 V, A 52 Yes Yes Yes Yes MC & OE
KnowIT-VQA [19] TV shows M 60 12 V, A, S 24 Yes No Yes Yes MC
AGQA [20] Social Media A 30 9.6 V, A 192 Yes No Yes Yes OE
AVQA [21] Social Media M 60 9.3 V, A 57.3 Yes No Yes Yes MC
STAR [22] Daily life A - 22 V 60 Yes No No Yes MC
Env-QA [23] Egocentric Video M 20 23.3 V 85.1 Yes No No Yes MC
FIBER [24] Daily life M 10 28 V, A, S 2 No Yes No Yes OE
HumorQA Daily life M 7 1.8 V, A 141.3 Yes Yes Yes Yes OE & FT
CreativeQA Performance M 48 0.9 V, A 78.7 Yes Yes Yes Yes OE & FT
MagicQA Magic M 10 1.6 V, A 91.9 Yes Yes Yes Yes OE & FT
FunQA Surprising videos M 19 4.3 V, A 312 Yes Yes Yes Yes OE & FT
@@
Video Question Answering Benchmarks While the visual question answering (VQA) task focuses on enhancing models’ ability in image comprehension [25, 26, 27], video question answering (VideoQA) shifts the attention towards video comprehension. VideoQA is generally more challenging than VQA as it requires a comprehensive understanding of visual content, utilization of temporal and spatial information, and exploration of relationships between recognized objects and activities [14]. To address the VideoQA task, the research community has introduced various benchmarks. As depicted in Table 1, Most commonly used VideoQA datasets are sourced from human-centric videos like movies [13], TV shows [15, 16, 19], and social media [12, 17, 18, 20, 21, 22, 24], and there are also object-centric datasets of game videos [11], synthetic videos [14] and egocentric videos [23]. MovieQA [13] and TVQA [15] are commonly employed by VideoQA methods, which put forward tasks related to temporal and causal reasoning. However, they rely heavily on dialogue comprehension and textual plot summaries, which severely limits the challenge of visual reasoning. TGIF-QA [12] uses animated GIFs to challenge spatial-temporal reasoning, but as most GIFs are short videos of 3 seconds, and its tasks mainly focus on action description, TGIF-QA lacks complex reasoning evaluation ability. When most datasets use multiple choice questions as QA tasks, some methods, such as NExT-QA [18], try to join open-ended questions. NExT-QA mainly focuses on daily life videos, but the open-ended answers are mostly simple sentences containing only a few words. To sum up, most existing methods focus on ordinary videos, lack of understanding of intriguing or unexpected videos, and advanced reasoning tasks such as generating complete explanatory texts of videos remain to be explored.
@@
Video Question Answering Solutions Earlier studies have explored various models, including LSTMs and graph-based neural networks, to capture cross-modal information [28, 29]. With the advent of Transformers, video understanding models like ClipBERT [30] and CoMVT [31] emerged, focusing on the comprehension of specific frames within a video. Subsequent models, such as Violet [32], extended their ability to encompass temporal and spatial information. However, these methods have primarily been applied to short videos. In the realm of long videos, MIST [33] stands out by achieving state-of-the-art performance and excelling in terms of computation efficiency and interpretability. Furthermore, recent Vision Language Models (VLMs) such as [34, 35, 36] have showcased remarkable video understanding capabilities.
@@
Counter-Intuitive Benchmarks While many current computer vision benchmarks primarily focus on understanding commonsense content, there is a growing interest in addressing the realm of counter-intuitiveness. Several emerging benchmarks and models cater to this domain, such as Whoops [37], which emphasizes weird, unusual, and uncanny images, and MemeGraphs [38], which revolves around memes featuring humor and sarcasm. Furthermore, some work even challenges models to comprehend complex multimodal humor in comics [39]. In the realm of large vision-language models, exemplified by GPT-4 [40], there is a particular focus on showcasing their ability to provide explanations for funny pictures. However, when it comes to videos, existing datasets exploring humor in TV shows or comedy tend to heavily rely on audio and narrative cues [8, 9, 10], with visual clues playing a comparatively lesser role.
3 FunQA Dataset
In this section, we provide a detailed explanation of the design principles that guided the creation of the FunQA dataset and its subsets. Additionally, we introduce our novel VideoQA tasks tailored for FunQA, and FunQA data statistics in Figure 2. We introduce our construction pipeline in the end.
(a) Word cloud of HumorQA
(b) Word cloud of CreativeQA
(c) Word cloud of MagicQA
(d) Number of videos and split
(e) Total video length
(f) Number of QA pairs
(g) Time Dist. of localization task
(h) Avg. word count of answers
(i) Anno. consensus evaluation
@@ 3 FunQA Dataset
Figure 2: Statistics of FunQA Dataset. FunQA consists of three subsets, each corresponding to different video types, and is annotated with free-text QA pairs. The first row displays word clouds representing critical annotations for each subset. The second row provides key dataset statistics, including the number of videos for different splits, video length, and QA pair count for three subsets. In the last row, (g) highlights the high-frequency time span of the answer for localization questions in red, (h) shows the average word count of answers, and (i) presents the percentage of consensus between annotators for the same answer in a sampled set.
3.1 Video Selection

### v2

@@ Abstract
Surprising videos, e.g., funny clips, creative performances, or visual illusions, attract significant attention. Enjoyment of these videos is not simply a response to visual stimuli; rather, it hinges on the human capacity to understand (and appreciate) commonsense violations depicted in these videos. We introduce FunQA, a challenging video question answering (QA) dataset specifically designed to evaluate and enhance the depth of video reasoning based on counter-intuitive and fun videos. Unlike most video QA benchmarks which focus on less surprising contexts, e.g., cooking or instructional videos, FunQA covers three previously unexplored types of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. For each subset, we establish rigorous QA tasks designed to assess the model’s capability in counter-intuitive timestamp localization, detailed video description, and reasoning around counter-intuitiveness. We also pose higher-level tasks, such as attributing a fitting and vivid title to the video, and scoring the video creativity. In total, the FunQA benchmark consists of 312K free-text QA pairs derived from 4.3K video clips, spanning a total of 24 video hours. Moreover, we propose FunMentor, an agent designed for Vision-Language Models (VLMs) that uses multi-turn dialogues to enhance models’ understanding of counter-intuitiveness. Extensive experiments with existing VLMs demonstrate the effectiveness of FunMentor and reveal significant performance gaps for the FunQA videos across spatial-temporal reasoning, visual-centered reasoning, and free-text generation.
@@ 1 Introduction
The charm of surprising videos, being funny, creative, and filled with visual illusions, offers enjoyment and attracts engagement from viewers. This type of media elicits positive surprise111c.f., negative surprise, e.g., a surprising medical bill. [46], a captivating emotion that stems not merely from perceiving surface-level visual stimuli, but rather, the innate ability of humans to understand and find delight in unexpected and counter-intuitive moments [40]. However, despite significant advancements in today’s computer vision models, the question remains: can video models “understand” the humor/creativity in surprising videos? Consider the humorous video depicted in Fig. 1 (left) as an example. We observe a woman in black holding two pot lids and clapping them together. The remaining three individuals are responsible for avoiding the pot lids. The first two people successfully dodge, but the third girl, in a panic, fails to avoid any hits and gets struck three times. The embarrassed demeanor of the third girl along with her final frustrated reaction, elicits laughter [^2] The hostility/superiority theory of humor posits that humor can arise from claiming superiority over someone or something [17, 2]; but alternate (more optimistic) theories of humor exist, [1] offers a survey.. While humans effortlessly recognize this as an unusual (and potentially entertaining) event, the reasoning required to holistically understand the scene is complex: a model needs to recognize that this is not a video depicting harm but rather girls engaging in playful pranks together, and discern that the comedic element arises from the stark contrast between the third girl being hit by the pot lids every time and the first two girls skillfully avoiding them.
Table 1: Comparison between FunQA and other existing benchmarks. Compared to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. For annotation type: [Uncaptioned image] denotes Manual Annotation and [Uncaptioned image] for Automatic Annotation; Avg Len denotes video average length; # Clips means number of video clips; VC for visual-centric, Des. for Description, Exp. for Explanation, STR for Spatial-temporal Reasoning, MC means Multiple Choice QA, and OE shows Open Ended QA with Average Word Count per response.
Dataset Domain [Uncaptioned image] or [Uncaptioned image] Video Question Answer
Avg Len # Clips # QA VC Des. Exp. STR MC OE
TGIF-QA [20] Social Media [Uncaptioned image] 3s 72K 165K 2.1
MSRVTT-QA [60] Social Media [Uncaptioned image] 15s 10K 244K 1.0
ActivityNet-QA [68] Social Media [Uncaptioned image] 180s 6K 58K 1.9
NExT-QA [59] Daily life [Uncaptioned image] 44s 5K 52K 2.6
Social-IQ [69] Daily life [Uncaptioned image] 99s 1K 8K N/A
MovieQA [54] TV shows [Uncaptioned image] 203s 7K 6K N/A
TVQA+ [33] TV shows [Uncaptioned image] 8s 4K 30K N/A
SUTD-TrafficQA [61] Traffic [Uncaptioned image] 5s 10K 623K N/A
MarioQA [44] Games [Uncaptioned image] 5s 188K 188K 2.0
CLEVRER [67] Synthetic Videos [Uncaptioned image] 5s 20K 305K N/A
FunQA (Ours) Surprising Videos [Uncaptioned image] 19s 4K 312K 34.2
@@
While there have been some efforts to enhance computer vision models’ performance in Video Question Answering (VideoQA), these works have primarily focused on the common, less surprising videos found in existing VideoQA datasets. Examples of commonly employed VideoQA datasets include YouCook2 [72] which contains video clips from 2K cooking videos, Howto100M [41] which consists of only instructional videos. While there exist video datasets that explore the humor in TV shows [5, 18] and include tasks such as predicting laughter tracks [49], these tasks often heavily rely on audio and narrative cues, with visual clues might playing a lesser role. Beyond datasets centered on factual queries, it is worth noting that NExT-QA targets the explanation of video content, which is widely employed for evaluating reasoning abilities. However, it was found in the experiment (see Section LABEL:sec:compare_nextqa) that VLMs such as GPT-4V(ision) already achieved an accuracy of 80% on NExTQA. This demonstrates that with the development of VLMs, the demand for datasets with deeper reasoning capabilities and presenting greater challenges is increasing.
@@
To revitalize the visual reasoning field and further improve model capabilities to identify and understand visual commonsense violations in videos, we introduce FunQA, an extensive and carefully curated VideoQA dataset comprising 4.3K surprising videos and 312K manually annotated free-text QA pairs. Unlike some VideoQA datasets that feature open-ended questions but short answers (e.g., an average of 2.6 words per answer in NExT-QA [59]), FunQA’s responses average 34.2 words in length. This significantly increases the demand for advanced video comprehension capabilities in the model. Therefore, here we use free-text QA to distinguish from open-ended QA. Our dataset consists of three subsets: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset covers different sources and contents, but the commonality lies in their surprising nature, e.g., the unexpected contrasts in humorous videos, the intriguing disguises in creative videos, and the seemingly impossible performances in magic videos. Our experiments suggest that these surprising videos require different types of reasoning than common videos, as existing VideoQA methods perform poorly on the corpus. With FunQA, we hope to provide a benchmark that covers the popular, important, and sophisticated genre of counter-intuitive/surprising videos.
@@
In FunQA, we formulate three rigorous tasks to measure models’ understanding of surprise: 1) Counter-intuitive timestamp localization: a model must identify the specific time period within a video when an unexpected event takes place. 2) Detailed video description: a model must generate coherent and objective descriptions of the video content, evaluating models’ fundamental video understanding capabilities. 3) Counter-intuitiveness reasoning: a model must generate concrete explanations of why the video is surprising. These tasks progressively assess the model’s ability to perceive, articulate, and reason about the counter-intuitive elements present in surprising videos. We also propose additional tasks that pose higher-level challenges, such as assigning an appropriate and vivid title to the video.
In continuation of our efforts to enhance models’ comprehension of surprising content, we introduce FunMentor, a specialized agent designed to boost counter-intuitive reasoning in VLMs. Operating like a seasoned coach in a variety show, FunMentor engages in detailed, multi-turn dialogues, honing the models’ responses to accurately grasp the essence of both amusing and astonishing content. Beyond just evaluation, FunMentor actively steers VLMs with precise prompts, fostering fluent, logical, and persuasive responses. Experiments have demonstrated its effectiveness in augmenting VLMs’ ability to comprehend.
To summarize our contributions:
@@
2) Novel and Challenging Tasks: We design a number of novel tasks that allow the model to explore previously untouched problems, such as timestamp localization, and reasoning around counter-intuitiveness. These tasks push video reasoning beyond superficial descriptions, demanding deeper understanding.
3) Novel Method named FunMentor: We propose a novel agent that refines the model’s understanding of counter-intuitiveness through multi-turn dialogues with VLMs.
4) Comprehensive Evaluation: We have done an extensive and comprehensive evaluation of cutting-edge baselines, giving the field an insight and future research direction.
@@ 2 Related Work
Video Question Answering Benchmarks While the visual question answering (VQA) task focuses on enhancing models’ ability in image comprehension [15, 73, 28], video question answering (VideoQA) shifts the attention towards video comprehension. VideoQA is generally more challenging than VQA as it requires a comprehensive understanding of visual content, utilization of temporal and spatial information, and exploration of relationships between recognized objects and activities [67]. To address the VideoQA task, the research community has introduced various benchmarks. As depicted in Table 1 (the complete table are shown in Appendix 0.A), most commonly used VideoQA datasets are sourced from human-centric videos like movies [54], TV shows [32, 33, 14], and social media [20, 69, 59, 16, 64, 58, 6], and there are also object-centric datasets of game videos [44], synthetic videos [67] and egocentric videos [12]. MovieQA [54] and TVQA [32] are commonly employed by VideoQA methods, which put forward tasks related to temporal and causal reasoning. However, they rely heavily on dialogue comprehension and textual plot summaries, which severely limits the challenge of visual reasoning. TGIF-QA [20] uses animated GIFs to challenge spatial-temporal reasoning, but as most GIFs are short videos of 3 seconds, and its tasks mainly focus on action description, TGIF-QA lacks complex reasoning evaluation ability. When most datasets use multiple choice questions as QA tasks, some methods, such as NExT-QA [59], try to join open-ended questions. NExT-QA mainly focuses on daily life videos, but the open-ended answers are mostly simple sentences containing only a few words. To sum up, most existing methods focus on ordinary videos, lack of understanding of intriguing or unexpected videos, and advanced reasoning tasks such as generating complete explanatory texts of videos remain to be explored.
@@
Video Question Answering Methods Earlier studies have explored various models, including LSTMs and graph-based neural networks, to capture cross-modal information [37, 71]. With the advent of Transformers, video understanding models, like ClipBERT [31] and CoMVT [53] emerged, focusing on understanding specific frames within a video. Subsequent models like Violet [11], extended their ability to encompass temporal and spatial information. However, these methods have primarily been applied to short videos. For long videos, MIST [13] stands out by achieving state-of-the-art (SOTA) performance and excelling in terms of computation efficiency and interpretability. Furthermore, recent VLMs [34, 36, 43] have showcased remarkable video understanding capabilities.
@@
Counter-Intuitive Benchmarks While many current computer vision benchmarks primarily focus on understanding commonsense content, there is a growing interest in addressing the realm of counter-intuitiveness. Several emerging benchmarks and models cater to this domain, such as Whoops [3], which emphasizes weird, unusual, and uncanny images, OOPS [9], which centers on recognizing and predicting unintentional events, and MemeGraphs [26], which revolves around memes featuring humor and sarcasm. Furthermore, some works even challenges models to comprehend complex multimodal humor in comics [19]. In the realm of large language models, exemplified by GPT-4 [47], there is a particular focus on showcasing their ability to provide explanations for funny pictures. However, regarding videos, existing datasets exploring humor in TV shows or comedy tend to heavily rely on audio and narrative cues  [5, 18, 49], with visual clues playing a comparatively lesser role.
3 The FunQA Dataset
(a) Number of videos and splits
(b) Number of QA pairs
(c) Time Dist. of Localization
(d) Consensus Evaluation
@@ 3 The FunQA Dataset
Figure 2: Statistics of FunQA Dataset. Figure (a) and (b) showcase vital statistics, including the number of videos for different videos, splits, and QA pairs count for three subsets. Figure (c) highlights the high-frequency time span of the answer for localization questions in red. Figure (d) presents the percentage of consensus between annotators for the same QA pair. The consensus is categorized into three levels High for consistent understanding, Medium for partial agreement but with mutual acknowledgment, and Low for complete disagreement.
In this section, we provide a detailed explanation of the design principles that guided the creation of the FunQA dataset and its subsets. Additionally, we introduce our novel VideoQA tasks tailored for FunQA, and FunQA data statistics in Figure 2. Toward the end we present our Construction Pipeline and Quality Control, highlighting our efforts in maintaining data quality and objectivity.
3.1 Task Definition
Detailed Description Task The description task aims to evaluate the model’s information extraction capabilities, serving as a fundamental aspect of video understanding. This task requires providing a free-text answer that describes the selected moment. Furthermore, this task allows for analysis of how the model extracts information and generates answers for subsequent tasks. By examining the model’s performance in this task, we gain insights into its ability to extract relevant information and generate meaningful responses.
Higher Level Tasks In addition to the three main tasks, we design higher-level tasks to enhance the model’s inference abilities on counter-intuitive videos. Title Task in HumorQA and CreativeQA requires generating a concise title summarizing the video’s content. Creative Scoring Task in CreativeQA involves rating the creativity of videos between 1 and 20 (provided officially). Magic Method Task in MagicQA requires the model to explain clearly the rationale behind the magic, and its purpose is to test the model’s ability to reason more deeply. To ensure the accuracy of the answers, this task is only partially annotated and appears only in the test set, details of which can be found in Appendix 0.A.1.
3.2 Dataset Statistics
FunQA contains 4,365 counter-intuitive video clips and 311,950 QA pairs, the total length of these videos is 23.9h and the average length of each clips is 19 seconds. FunQA consists three fine-grained subsets, each one containing well-designed tasks. The specific numbers of videos and splits can be seen in Fig. 2 (a). The specific number of QA pairs for each task can be seen in Fig. 2 (b). For our localization task, the heat map for the three different types of videos can be seen in Fig. 2 (c), which shows the high-frequency time span of the answer. For the description and reasoning tasks, the average length of the words in their free-text answers reached 34.24, which is much longer than existing VideoQA datasets (e.g., 2.6 in NExT-QA [59]). FunQA has a well-established annotation process and high annotation quality, the result of our annotation consensus evaluation are illustrated in Fig. 2 (d). Impressively, over 90% of the annotations demonstrate a high level of consensus, while only 1% exhibit low consensus. This clearly underscores the objectivity and reliability of the FunQA dataset.
3.3 Dataset Construction Pipeline
FunQA dataset construction pipeline was in three stages: Pre-processing, Manual Annotation, and Post-Processing. The whole process took about 900 hours with over 50 highly educated undergraduates as part-time annotators. See Appendix 0.A.1 for more details on the dataset construction.
Pre-Processing Initially, we crawled videos from YouTube. Then we performed a two-stage manual cleaning and cutting process on the collection to ensure counter-intuitive features and video quality and to exclude non-ethical and sensitive content, resulting in video clips.
Manual Annotation We annotated the videos according to the characteristics of different task designs in Chinese. We screen and train the annotators to ensure the accuracy and high quality of the annotation, and finally produce the original annotated files. After the first round, we conducted a secondary round of 10% of the tasks and performed Consensus Evaluation to ensure the objectivity.
Post-Processing Based on our carefully designed tasks and high-quality annotations, we expanded our dataset using GPT-3.5. Firstly, we automatically translated the Chinese annotations into English. Subsequently, we generated more QA pairs that were faithful to the original ideas but presented differently. This not only made FunQA multilingual but also expanded its QA pair count to 312K. Additionally, we created diverse sub-dataset, FunQA-MC (multi-choice QA) and FunQA-DIA (dialogue QA). In addition, to focus on exploring the ability to handle counter-intuitive reasoning, we released FunQA-MC-R (a multi-choice version specifically containing counter-intuitive reasoning questions). More details are given in Appendix 0.A.2 and Appendix 0.B.
Table 2: Consensus Evaluation Experiment. This table shows the results from a random 10% sample of QA pairs, cross-validated by annotators to assess agreement with existing annotations. ‘Low,’ ‘Medium,’ and ‘High’ indicate the strength of the consensus. Consensus HumorQA CreativeQA MagicQA # Total Low 1 1 2 4 (1%) Medium 9 6 17 32 (8%) High 199 111 194 504 (91%) Table 3: Can FunQA be Solved Solely Based on Images? The left two columns show the average number of questions answerable or not by humans using only 8 static, uniformly selected frames. Dataset # Can # Cannot Cannot Rate HumorQA 7.8 32.2 80% CreativeQA 6.1 33.9 85% MagicQA 2.6 37.4 94% FunQA 16.5 103.5 86%
3.4 Quality Control
Minimal Errors and High Objectivity in FunQA We assure that every annotation included in the final release of FunQA has been subjected to rigorous multi-person, multi-round review processes. Furthermore, We did manual consensus evaluation on released FunQA dataset, randomly sampling 10% of the data from all three sub-datasets (HumorQA, CreativeQA, and MagicQA). As shown in Table 3.3, we get the 91% high consensus.
FunQA Emphasis on Temporal Dynamics FunQA requires a strong emphases on temporal dynamics rather than solely on few frames of images. To prove that, we did the quantitative human experiments - We randomly selected 40 videos from each of the three sub-datasets, totaling 120 videos. For each video, we sampled 8 frames evenly. We enlisted 10 individuals who had not seen any FunQA videos before. We had them view the sequence of 8 consecutive frames and then watch the original video along with its annotations. They were asked to determine whether they can understand and answer the counter-intuitive understanding of the original video solely based on the images. Nearly 86% people thought that FunQA cannot be solved only by images, as shown in Table 3.3.
4 Limitations and Future Work
This paper has two limitations. 1) The current FunQA dataset primarily contains video-level data and annotations. There is potential for enhanced video reasoning through denser annotations, akin to PVSG [63], which might include detailed spatial, temporal, and object-level annotations. 2) The initial annotations were made in Chinese and later translated into English. While GPT was used to refine and complete the Chinese text, ensuring comprehensiveness and correctness, differences due to cultural differences between the languages might still persist.
Looking ahead, we plan to enrich FunQA with more detailed and varied annotations. We aim to develop new metrics for a more accurate evaluation of models, particularly for open-ended questions. Our goal is to steer models towards deeper video reasoning. However, to ensure fair comparisons and prevent data leakage, it is advised that future research does not utilize the FunQA testing set.
5 Acknowledgments and Disclosure of Funding
Appendix 0.A More Details of FunQA Dataset
0.A.1 Dataset Construction Pipeline